Solve the differential equation of the following homogeneous equation.

xydx + (x^2+ y^2)dy = 0
1
Expert's answer
2021-09-19T18:06:53-0400

Let us solve the differential equation "xydx + (x^2+ y^2)dy = 0." For this let us use the transformation "y=ux." Then "dy=udx+xdu," and hence we get the differential equation "ux^2dx + (x^2+ u^2x^2)(udx+xdu) = 0." It follows that "ux^2dx + x^2(1+u^2)udx+ (1+ u^2)x^3du= 0," and hence "x^2(2u+u^3)dx+ (1+ u^2)x^3du= 0."

Consequently, we have the equation "\frac{dx}{x}+ \frac{(1+ u^2)du}{u(2+u^2)}= 0." It follows that "\int\frac{dx}{x}+ \int\frac{(1+ u^2)du}{u(2+u^2)}= \ln|C|."

We conclude that "\ln|x|+ \frac{1}{2}\int(\frac{1}{u}+\frac{u}{2+u^2})du= \ln|C|," and hence "\ln|x|+ \frac{1}{2}\ln|u|+\frac{1}{4}\ln(2+u^2)= \ln|C|." Therefore, "4\ln|x|+ 2\ln|u|+\ln(2+u^2)= \ln|C_1|."

We conclude "\ln(x^4u^2(2+u^2))= \ln|C_1|," and thus "x^4u^2(2+u^2)=C_1." Then "x^4(\frac{y}x)^2(2+(\frac{y}x)^2)=C_1." We conclude that the general solution of the equation is "y^2(2x^2+y^2)=C."